OpenStudy (anonymous):

Solve the equation: (√x-1)=x-7

OpenStudy (shamil98):

(√x-1) = x - 7 square both sides x-1 = (x-7)^2 can you do it from here?

OpenStudy (anonymous):

When you square the square root it turns into an absolute value, right?

OpenStudy (shamil98):

when you square a number in a square root symbol \[\huge (\sqrt{x})^2 = x\] you get that number inside the square root symbol

OpenStudy (anonymous):

and I would factor out the (x-7)^2?

OpenStudy (shamil98):

yes, multiply (x-7)(x-7)

OpenStudy (anonymous):

Now I have: x-1=x^2-14x+49, am I suppose to add the 1 to both sides?

OpenStudy (anonymous):

You arrange it so that 0= something. So you add one to both sides and also subract x from both sides.

OpenStudy (anonymous):

You were right to be wary of the squaring operation though. You will get 2 solutions when you factor, only one of them actually satisfies the original equation.

OpenStudy (anonymous):

Once I minus the x from both sides and add the one, it looks like: 0=x^2-13x+50, right?

OpenStudy (anonymous):

-14x -x=-15x

OpenStudy (anonymous):

0=x^2-15x+50

OpenStudy (anonymous):

Yeah that's right, I didn't notice that I did that.

OpenStudy (anonymous):

So now we have to factor that. :)

OpenStudy (anonymous):

How come the x=5 doesn't satisfy the equation?

OpenStudy (anonymous):

Because of the squaring we did to get rid of the radical.

OpenStudy (anonymous):

The easiest way to solve these though is just how we did it, and then know that you have to check your answers to see which actually work.

OpenStudy (anonymous):

Oh ok I see what you mean. Thank you for your help!

OpenStudy (anonymous):

No problem :)